OpenStudy (anonymous):

an angle measures 62 degrees. what is the measure of its complement

OpenStudy (anonymous):

the complement of an angle is the angle measure that when added together would resutl in 90 degrees. in this case: 62 + x = 90 x = 90 - 62 x = 28

OpenStudy (anonymous):

The difference between 90 degrees and an angle is its complement. 90 - 62 = 28 degrees.